OpenStudy (anonymous):

. Having survived the meteor impact, thanks to some last minute evasive maneuvers, the Mathonauts now set their sights on their Interstellar Headquarters. The Interstellar Headquarters orbits the Earth based on the equation y2 + x2 = 40,000. Using the original trajectory of the ship and complete sentences, explain to the pilot how to find where the ship’s path will cross the Interstellar Headquarters’s path.

OpenStudy (anonymous):

My original equation is y=2x+4

OpenStudy (anonymous):

What I thought would be the first step in this process is that I first square root the equation y2 + x2 = 40,000. correct @mathstudent55

OpenStudy (mathstudent55):

You can't take the square root of the second equation. You need to solve the two equations simultaneously.

OpenStudy (mathstudent55):

You can use the substitution method. Since the equation y = 2x + 4 is already solved for y, substitute y in the other equation with y^2 + x^2 = 40,000.

OpenStudy (anonymous):

Ok what that would mean is that I plug in the y value I had gotten which is 2x+4 and square that which would cause for me to foil

OpenStudy (mathstudent55):

Yes.

OpenStudy (mathstudent55):

\(y = 2x + 4\) \(y^2 + x^2 = 40000\) Substitute y in the second equation with \(2x + 4\). \((2x + 4)^2 + x^2 = 40000\)

OpenStudy (anonymous):

If I foiled it I got 4x^2+16x^2+16

OpenStudy (anonymous):

Which if I simplified again i GOT 18X^2

OpenStudy (anonymous):

Which the equation I got was 18x^2+16

OpenStudy (anonymous):

@mathstudent55

OpenStudy (mathstudent55):

This is how you square a binomial: \((a + b)^2 = a^2 + 2ab + b^2\) In your case, you get: \((2x + 4)^2 + x^2 = 40000\) \(4x^2 + 16x + 16 + x^2 = 40000\)

OpenStudy (mathstudent55):

Using FOIL you got it almost right. The 16x term is just 16x, not 16x^2.

OpenStudy (mathstudent55):

Do you understand it so far?

OpenStudy (anonymous):

Yes

OpenStudy (mathstudent55):

Using FOIL, squaring your binomial is like this: \((2x + 4)(2x + 4) = 4x^2 + 8x + 8x + 16 = 4x^2 + 16x + 16\)

OpenStudy (mathstudent55):

Ok. Now we take the equation, and we combine like terms.

OpenStudy (mathstudent55):

\(4x^2 + 16x + 16 + x^2 = 40000\) 4x^2 and x^2 are like terms. Also, we subtract 40000 from both sides. \(4x^2 + x^2 + 16x + 16 - 40000 = 0\) \(5x^2 + 16x - 39984 = 0\)

OpenStudy (mathstudent55):

Now we need to solve that quadratic equation for x.

OpenStudy (anonymous):

ok

OpenStudy (mathstudent55):

We need to use the quadratic formula. Are you familiar with it?

OpenStudy (anonymous):

yes I am

OpenStudy (anonymous):

-b^2+squareroot of b^2-4ac/2a -

OpenStudy (mathstudent55):

The solution of the quadratic equation \(ax^2 + bxc + c = 0\) is \(x=\dfrac{-b \pm \sqrt{b^2 - 4ac}}{2a} \)

OpenStudy (mathstudent55):

Be careful. The first term is simply b, not b^2.

OpenStudy (anonymous):

yeah

OpenStudy (mathstudent55):

\(x = \dfrac{-16 \pm \sqrt{16^2 - 4(5)(-39984)}}{2(5)}\)

OpenStudy (anonymous):

I got -16+ squaroot of 79993.6 -

OpenStudy (mathstudent55):

\(x = \dfrac{-16 \pm \sqrt{256 + 799680}}{10}\) \(x = \dfrac{-16 \pm \sqrt{799936}}{10}\) \(x = \dfrac{-16 \pm \sqrt{799936}}{10}\)

OpenStudy (mathstudent55):

Now we need to simplify the root. \(x = \dfrac{-16 \pm \sqrt{64 \times 12499}}{10}\)

OpenStudy (mathstudent55):

\(x = \dfrac{-16 \pm 8\sqrt{12499}}{10}\)

OpenStudy (mathstudent55):

\(x = \dfrac{-8 \pm 4\sqrt{12499}}{5}\) \(x = \dfrac{-8 + 4\sqrt{12499}}{5}\) or \(x = \dfrac{-8 - 4\sqrt{12499}}{5}\)

OpenStudy (mathstudent55):

Now we plug in each x-value we got in the first equation, and we find each corresponding y-value.

OpenStudy (mathstudent55):

\(x = \dfrac{-8 + 4\sqrt{12499}}{5}\) or \(x = \dfrac{-8 - 4\sqrt{12499}}{5}\) \(y = 2x + 4\) \(y = 2\times \dfrac{-8 + 4\sqrt{12499}}{5} + 4\); \(y = 2\times \dfrac{-8 - 4\sqrt{12499}}{5} + 4\) \(y = \dfrac{-16 + 8\sqrt{12499}}{5} + \dfrac{20}{5}\); \(y = \dfrac{-16 - 8\sqrt{12499}}{5} + \dfrac{20}{5}\) \(y = \dfrac{4 + 8\sqrt{12499}}{5}\); \(y = \dfrac{4 - 8\sqrt{12499}}{5}\) The solutions are: \(x = \dfrac{-8 + 4\sqrt{12499}}{5}\), \(y = \dfrac{4 + 8\sqrt{12499}}{5}\) \(x = \dfrac{-8 - 4\sqrt{12499}}{5}\), \(y = \dfrac{4 - 8\sqrt{12499}}{5}\)

OpenStudy (anonymous):

Thank you